Becton Labs, Inc., produces various chemical compounds for industrial use. One compound, called Fludex, is prepared using an elaborate distilling process. The company has developed standard costs for one unit of Fludex, as follows:

Standard Quantity
or Hours
Standard Price
or Rate
Standard Cost
Direct materials 2.40 ounces $ 21.00 per ounce $ 50.40
Direct labor 0.80 hours $ 15.00 per hour 12.00
Variable manufacturing overhead 0.80 hours $ 3.50 per hour 2.80
Total standard cost per unit $ 65.20

During November, the following activity was recorded related to the production of Fludex:

  1. Materials purchased, 13,500 ounces at a cost of $266,625.
  2. There was no beginning inventory of materials; however, at the end of the month, 4,200 ounces of material remained in ending inventory.

  3. The company employs 24 lab technicians to work on the production of Fludex. During November, they each worked an average of 140 hours at an average pay rate of $14.50 per hour.

  4. Variable manufacturing overhead is assigned to Fludex on the basis of direct labor-hours. Variable manufacturing overhead costs during November totaled $6,800.

  5. During November, the company produced 3,800 units of Fludex.

Required:

1. For direct materials:

a. Compute the price and quantity variances.

b. The materials were purchased from a new supplier who is anxious to enter into a long-term purchase contract. Would you recommend that the company sign the contract?

2. For direct labor:

a. Compute the rate and efficiency variances.

b. In the past, the 24 technicians employed in the production of Fludex consisted of 4 senior technicians and 20 assistants. During November, the company experimented with fewer senior technicians and more assistants in order to reduce labor costs. Would you recommend that the new labor mix be continued?

3. Compute the variable overhead rate and efficiency variances.
Answer #1

1.
Material Price Variance = (Actual Price - Standard Price) x Actual Quantity purchased
= ($19.75 - $21) x 13500 = $16875 (F)
Actual price = $266625 / 13500 = $19.75 per ounce

Material Quantity Variance = (Actual Quantity used - Standard Quantity) x Standard price
= (9300 - 9120) x $21 = $3780 (U)
Actual Quantity used = 13500 - 4200 = 9300
Standard Quantity = 3800 x 2.40 = 9120

b.
Even though there is unfavorable material quantity variance, but material price variance is favorable by higher amount.
So yes it is recommended to sign the contract.

2.
Direct Labor rate variance = (Actual rate - Standard rate) x Actual hours
= ($14.50 - $15) x 3360 = $1680 (F)
Actual hours = 140 x 24 = 3360

Direct Labor efficiency variance = (Actual hours - Standard hours) x standard rate
= (3360 - 3040) x $15 = $4800 (U)
Standard hours = 3800 x 0.80 = 3040

b. Not recommended
Due to higher unfavorable direct labor rate variance

3.
Variable overhead rate variance = (Actual rate - Standard rate) x Actual hours
= $6800 - $3.50 x 3360 = $4960 (F)
Actual hours = 140 x 24 = 3360

Variable overhead efficiency variance = (Actual hours - Standard hours) x standard rate
= (3360 - 3040) x $3.50 = $1120 (U)
Standard hours = 3800 x 0.80 = 3040